OpenStudy (mrdoodbromanguy):

. Jennifer chooses a playing card at random from a deck of 52 cards then she chooses another one. this one i'm unsure of but i think it has something to do with probability

OpenStudy (tt0860):

There is a \frac{13}{52} chance she will pick a heart the first time and a \frac{12}{51} that she will the second time. The probability of this occurring thus is: \frac{13}{52}* \frac{12}{51} = \frac{13*2*2*3}{2*2*13*3*17} = \frac{1}{17}

zepdrix (zepdrix):

Is Jennifer trying to choose hearts consecutively? Hmm I don't see that in the question...

Directrix (directrix):

@mrdoodbromanguy What are the instructions for this problem?